# Brain-Inspired Perspective on Configurations: Unsupervised Similarity and Early Cognition

### v1

@@ 3.1 Recap of Configuration Definition
Actually, the implementation of [2] (also the implementation of this paper), when $\gamma$ is fixed, is nothing big but the Leiden method [25] on $k$ nearest neighbor graph, with some minor modifications. As the graph is sparse with a small $k$, the runtime is essentially linear in $n$, i.e., $O(n)$. However, it can be exhaustive to search $\gamma\in[0,\infty)$ for “best” configurations. Luckily, [2] provides a solution called Parallel-DT. Parallel-DT is guaranteed to find $m+1$ segments of $(0,\infty)$ with dominant ${\bm{\Omega}}_{i}$ at the division points, and such segments are called plateaus. Combining all ${\bm{\Omega}}_{i}$, we get the desired $\bm{\Omega}$}.
@@ 5 Experiments
Infant Behavior Validation: Following [3], we test methods on internal cues (details like eyes, noses), external cues (outer shape of face/body), or both. This directly mirrors infant studies where 3-4-month-olds showed poor categorization with isolated cues but strong performance with combined cues. Table˜2 compares computational methods with infant behavior.

### v2

@@ 3.1 Recap of Configuration Definition
Actually, the implementation of [36] (also the implementation of this paper), when $\gamma$ is fixed, is nothing big but the Leiden method [43] on $k$ nearest neighbor graph, with some minor modifications. In specific, all edge weights are normalized so that the total weight of a column sums to one. Rather than fixing a single resolution, $\gamma$ is scanned dynamically and convergence is defined by configuration stability: cluster memberships remain unchanged under small perturbations of $\gamma$. This stability-based stopping replaces reliance on a single modularity peak and yields reproducible behavior and fair cross-dataset comparison. As the graph is sparse with a small $k$, the runtime is essentially linear in $n$, i.e., $O(n)$. However, it can be exhaustive to search $\gamma\in[0,\infty)$ for “best” configurations. Luckily, [36] provides a solution called Parallel-DT. Following [29, 36], Parallel-DT is a resolution-free procedure that discovers a chain of stable configurations between the one-cluster and all-singleton extremes. It iteratively identifies intermediate configurations that remain unchanged over ranges of $\gamma$ (configuration plateaus), running multiple local searches in parallel for efficiency, with a Leiden-based inner optimizer. Parallel-DT is guaranteed to find $m+1$ segments of $(0,\infty)$ with dominant ${\bm{\Omega}}_{i}$ at the division points, and such segments are called plateaus. Combining all ${\bm{\Omega}}_{i}$, we get the desired $\bm{\Omega}$}.
@@ 5 Experiments
Sensitivity and RMS ablation. RMS aligns labels across resolutions to avoid penalizing semantically coherent merge–split dynamics without changing the clustering itself. In Table˜2, Salinas shows substantial improvements (ARI +0.10, NMI +0.04), InfantS shows negligible changes (no greater than +0.01 on both), and ImageNet exhibits modest gains (ARI +0.03, NMI +0.03). These dataset-dependent effects match the expected merge–split sensitivity and support reproducible cross-resolution evaluation.
Table 2: RMS alignment ablation (ARI/NMI) by dataset.
Metric Salinas InfantS ImageNet
w/o RMS w/o RMS w/o RMS
ARI 0.82±0.01 0.92±0.01 0.55±0.01 0.55±0.02 0.59±0.02 0.62±0.02
NMI 0.90±0.01 0.94±0.01 0.57±0.02 0.58±0.02 0.65±0.02 0.68±0.01
Infant Behavior Validation: Following [39], we test methods on internal cues (details like eyes, noses), external cues (outer shape of face/body), or both. This directly mirrors infant studies where 3-4-month-olds showed poor categorization with isolated cues but strong performance with combined cues. Table˜3 compares computational methods with infant behavior.